### v1

@@ Abstract
To gain insight, small size and high-resolution hydrodynamical simulations have been performed (e.g. Greif et al., 2011; Wise et al., 2014; Stacy et al., 2016; Xu et al., 2016; Hirano et al., 2018; Wollenberg et al., 2020; Sugimura et al., 2020; Chon & Omukai, 2020; Chon et al., 2021, 2022; Toyouchi et al., 2023; Prole et al., 2023; Sadanari et al., 2024) that suggest that metal-free (or poor) mini-halos favour the formation of Pop. III stars with a more top-heavy Initial Mass Function (IMF) and with lower star formation efficiencies than observed today. It is also thought that Pop. III star formation might occur down to the end of the EoR at in pristine metal free pockets of gas (e.g. Venditti et al., 2023).
@@
The variety and complexity of the processes involved in Pop. III star formation and the resolution required to keep track of the evolution of the gas particles, limits the size of these hydrodynamical simulations to kpc. In order to mitigate this problem, semi-analytical models that account for Pop. III star formation have been developed (e.g. Visbal et al., 2020; Trinca et al., 2022; Hegde & Furlanetto, 2023; Feathers et al., 2024; Bovill et al., 2024; Ventura et al., 2024; Liu et al., 2024). These models allow a statistical study of Pop. III star formation in mini-halos out to scales of Mpc. While these volumes start to investigate the chemical enrichment of the intergalactic medium (IGM) and the Pop. III/II transition, they are still too small to study the EoR as volumes of at least Mpc are required (Iliev et al. 2014; Kaur et al. 2020; Balu et al. 2023a).
@@
Observations and models are converging on a scenario where the Universe was completely ionized by (e.g. Fan et al., 2006; Ouchi et al., 2010; McGreer et al., 2015; Chardin et al., 2015; Planck Collaboration et al., 2020; Qin et al., 2021a; Bosman et al., 2022; Jin et al., 2023; Qin et al., 2024) with reionization likely driven by low-mass halos (e.g. Kuhlen & Faucher-Giguère, 2012; Finkelstein et al., 2012; Robertson et al., 2015; Qin et al., 2021b; Saxena et al., 2024; Mutch et al., 2024). However, the impact of Pop. III stars and mini-halos on the EoR is unclear. Pop. III stars are likely to be the dominant contribution to the total star formation rate density (SFRD) at and, if their IMF is more top-heavy than the present day one, Pop. III could significantly contribute to the heating and the ionization of the IGM which determines the evolution and shape of the 21cm signal (e.g. Qin et al., 2020; Gessey-Jones et al., 2022; Sartorio et al., 2023).
@@
The 21cm signal represents our most promising tool to put constraints on the thermal state of the IGM during the Cosmic Dawn and EoR. Even though no confirmed detection has been reported so far, the first upper limits on the 21cm power spectrum obtained with HERA phase I strongly disfavour cold reionization scenarios (HERA Collaboration et al. 2023). In the last few years, the impact of Pop. III stars on the 21cm signal has been studied using both analytical and semi-analytical models (e.g. Cohen et al., 2017; Chatterjee et al., 2020; Mebane et al., 2020; Qin et al., 2020; Magg et al., 2022; Gessey-Jones et al., 2022; Muñoz et al., 2022; Hegde & Furlanetto, 2023; Ventura et al., 2023; Sartorio et al., 2023; Cruz et al., 2024a). However these models either did not compute reionization (e.g. Magg et al. 2022; Hegde & Furlanetto 2023; Cruz et al. 2024a), focusing only on the absorption trough of the 21cm signal occurring at , or used a very simple analytical approach to compute reionization (e.g.Ventura et al. 2023). On the other hand, Cohen et al. (2017), Qin et al. (2021a) and Muñoz et al. (2022) used a simple analytical model for modeling Pop. III star formation but computed the reionization self-consistently.
@@
In this work, we overcome these challenges using a realistic Pop. III star formation and mini-halo model (Ventura et al. 2024) developed within the semi-analytical model meraxes designed to self-consistently couple galaxy formation and reionization. While in Ventura et al. (2024) we ran this model on a small (L 10 Mpc) and high-resolution box, here we extend it to a significantly larger volume simulation (L cMpc) enabling the study of cosmic reionization. Since at such large volumes we cannot directly resolve mini-halos, we implemented scaling relations between the SFRD and the dark matter density field calibrated on the results from the small and high-resolution box discussed in Ventura et al. (2024). With this new model we are able to accurately follow the evolution of the radiative backgrounds relevant to the EoR and 21cm signal (X-rays, Lyman-, ionizing UV, Lyman-Werner) and to disentangle the contribution of Pop. III star formation to the 21cm global signal and power spectrum. Pop. III stars are expected to have a stronger impact at where they dominate star formation and ionization. In this work we focus our attention on the residual signature of Pop. III on the 21cm power spectrum at where the sensitivity of SKA is expected to be significantly better and a detection is more plausible. This study allows us to assess under which conditions an early heating of the IGM from Pop. III stars leaves a detectable imprint on the 21cm power spectrum at .
@@
This paper is structured as follows. In Section 2 we give a brief overview of Pop. III star formation in meraxes. In Section 3 we present the scaling relation between the SFRD in mini-halos and the density field calibrated from the small and high-resolution box which is implemented in the large (310 cMpc)3 box. In Section 4 we discuss the impact of different Pop. III star formation models on the 21cm power spectrum and in Section 5 we make forecasts on the observability of these power spectra with SKA. Finally, we summarize our main results and conclusions in Section 6. Our simulations use the best-fit parameters from the Planck Collaboration et al. (2016): h = 0.6751, = 0.3121, = 0.0490, = 0.6879, = 0.8150, and ns = 0.9653.
@@ 2.1 Galaxy formation
In V24 the cooling prescriptions were updated to account for H2 cooling (the main cooling channel in mini-halos) and a more detailed metal enrichment model to keep track of the metallicity of each gas reservoir in a halo (which is crucial to distinguish between Pop. III and Pop. II star formation episodes). We also account for the effects of both baryon-dark matter streaming velocity and H2 photo-dissociation by the Lyman-Werner background which increases the minimum mass of a mini-halo capable of hosting stars (e.g. Schauer et al., 2021). For this work, we updated meraxes by adding the effect of H2 self-shielding which counteracts the H2 photo-dissociation, increasing the Pop. III SFRD by up to one order of magnitude at (see e.g. Feathers et al. 2024). We discuss the details of the implementation and impact of H2 self-shielding in meraxes in the Appendix A.
@@
The main free parameters that regulate the galaxy formation in meraxes are summarized in Table 1. The Pop. II related ones are taken from Balu et al. (2023a) where meraxes was run on a cosmological volume of cMpc resolving all atomic cooling halos and calibrated in order to match the observed UV luminosity functions at (the agreement holds up to as shown in Qin et al. 2023) and the stellar mass functions at . Given the lack of observations of Pop. III stars, the Pop. III parameters are largely unconstrained. The fiducial values adopted in this work are taken from V24 and their values are suggested from hydrodynamical simulations (e.g. Chon et al., 2021).
@@ 2.2 Reionization and radiative backgrounds
where Tα is the colour temperature which we take equal to TK while xα and xc are the Lyman- and collisional coupling coefficients, respectively. These coefficients quantify the strength of the processes (resonant scattering of Lyman- photons, Wouthuysen 1952, and collisions with free electrons) that drive the spin temperature towards the kinetic temperature (when xα + xc 1, T TK otherwise TTγ). TK is sensitive to the adiabatic cooling and to all the processes able to heat up (or cool) the IGM with the most dominant coming from the X-ray emission. Hence in this work we will consider only the X-ray heating neglecting the other source of heating such as primordial magnetic fields (Minoda et al. 2019; Bera et al. 2020; Cruz et al. 2024b), Lyman- (Ciardi et al. 2010; Reis et al. 2021; Mittal & Kulkarni 2021), shocks (Furlanetto & Loeb 2004; Gnedin & Shaver 2004; Ma et al. 2021), cosmic rays (Bera et al. 2022), early accreting black holes (Mebane et al. 2020; Ventura et al. 2023) and decaying or annihilating dark matter (Liu & Slatyer 2018; Sun et al. 2023; Hou & Mack 2024; Facchinetti et al. 2024). Ultimately, the evolution of TS during the Cosmic Dawn and the EoR is mostly determined by the Lyman- (for the coupling between TS and TK) and X-ray flux.
@@
In this section, we present a novel approach that enables us to efficiently estimate the SFRD from mini-halos in a large box (that does not directly resolve these objects) using results from a small, high-resolution box.
@@ 3.1 Calibrating scaling relations from the L10 box
In Fig. 3 we show the redshift evolution of these parameters for 0.5 (black), 1.0 (grey), 1.5 (purple), 2.0 (red), 2.5 (green) and 3.0 (blue). exhibits an almost constant trend in redshift and a correlation with (higher results in higher ). This demonstrates that is mostly determined by the number of Pop. III star forming halos in a pixel, which is higher for more overdense regions. Since Pop. III star formation episodes in mini-halos are often the first episode of star formation experienced by a galaxy, it is not impacted by supernova feedback222Even though supernova feedback can be neglected, this is not true for the Lyman-Werner background that halts star formation in mini-halos. For this reason we add a further condition that if a pixel is irradiated by a LW flux J Jcrit above a critical threshold defined as M Mato in Eq. 8, (log10(SFR) = 0. so the Pop. III SFR is almost constant at all redshift.
@@ 3.2 Applying scaling relations to the L210 box
Differently from what has been done by Hazlett et al. (2024) who calibrated a semi-analytical model to the Reinassance simulation in order to account for Pop. III star formation by adding the previous star formation history to each atomic cooling galaxy resolved in the simulation, the methodology described in this section allows us to estimate only the total SFRD occurring in mini-halos within a certain pixel of the simulation. In this work we focus on three models of Pop. III star formation by varying three main parameters: the star formation efficiency, the IMF and the specific Pop. III X-ray luminosity per unit star formation while all the other free parameters (e.g. the escape fraction) are fixed at the fiducial value (see Tables 1 and 2). We chose to focus only on these three parameters since these have the strongest impact on both the evolution of the Pop. III SFRD and the amount of UV and X-ray photons emitted. The values chosen for the specific Pop. III X-ray luminosity per unit star formation are similar to those found in Sartorio et al. (2023) for the different IMFs explored in their work. Hereafter, we analyze three different Pop. III models designed to have the minimum, intermediate and maximum impact from Pop. III star formation in mini-halos, each model is summarized in Table 3.
@@
Extreme Pop. III | logE | 0.08 |
@@
We compute the sky-averaged 21cm global signal (see Fig. 10) without (black line) and with (grey, cyan and red line for weak, moderate and extreme Pop. III respectively) Pop. III star formation. As expected, introducing a Pop. III population with the same X-ray properties as Pop. II ones (i.e. weak Pop. III), simply shifts the absorption through to earlier epochs in virtue of the stronger coupling at higher-z (see also Ventura et al. 2023; Hegde & Furlanetto 2023). However, if Pop. III stars have a stronger X-ray emission (i.e. moderate and extreme models) as suggested by Sartorio et al. (2023), the absorption signal is quickly suppressed turning into an emission signal as early as for the extreme Pop. III model and at 13 for the moderate Pop. III one.
@@
As shown in Fig. 11 earlier coupling and heating from Pop. III impacts the 21cm power spectrum both at the large and small scales. The models considered here produce 21cm signals that are different not only during the coupling and heating epoch (), but also at lower redshift when the reionization is in progress. The impact is stronger at smaller scales where models with a stronger heating exhibit a larger power spectrum at . This is of crucial importance as current and upcoming facilities are improving their observations at z 10 (see more discussion in Section 5.
@@
This analysis shows that Pop. III star formation not only impacts the 21cm global signal during the cosmic dawn as previously assessed (e.g. Qin et al., 2020; Gessey-Jones et al., 2022; Ventura et al., 2023; Hegde & Furlanetto, 2023; Cruz et al., 2024a) but also that an early () heating of the IGM provided by this population leaves a strong signature on the power spectrum during the EoR (z ). The impact on the power spectrum is stronger for models that have a large Pop. III X-ray emissivity (i.e. moderate and extreme Pop. III models) while models with a milder X-ray emission (i.e. weak Pop. III) have a stronger effect on the global signal. Ultimately, this tells us that the power spectrum during the EoR can be used to disentangle different heating models and potentially constrain the properties of Pop. III stars. Finally, we highlight that here we considered only the X-ray emission from stellar remnants. There might be other sources that significantly heat the IGM at whose impact would add to the Pop. III one considered here.

### v2

@@ Abstract
To gain insight, small size and high-resolution hydrodynamical simulations have been performed (e.g. Greif et al., 2011; Hirano et al., 2018; Chon et al., 2021, 2022; Toyouchi et al., 2023; Sadanari et al., 2024) that suggest that metal-free (or poor) mini-halos favour the formation of Pop. III stars with a more top-heavy Initial Mass Function (IMF) and with lower star formation efficiencies than observed today. It is also thought that Pop. III star formation might occur down to the end of the EoR at in pristine metal free pockets of gas (e.g. Venditti et al., 2023).
@@
The variety and complexity of the processes involved in Pop. III star formation and the resolution required to keep track of the evolution of the gas particles, limits the size of these hydrodynamical simulations to kpc. In order to mitigate this problem, semi-analytical models that account for Pop. III star formation have been developed (e.g. Visbal et al., 2020; Hegde & Furlanetto, 2023; Liu et al., 2024). These models allow a statistical study of Pop. III star formation in mini-halos out to scales of Mpc. While these volumes start to investigate the chemical enrichment of the intergalactic medium (IGM) and the Pop. III/II transition, they are still too small to study the EoR as volumes of at least Mpc are required (Iliev et al. 2014; Kaur et al. 2020; Balu et al. 2023a).
@@
Observations and models are converging on a scenario where the Universe was completely ionized by (e.g. Fan et al., 2006; Ouchi et al., 2010; McGreer et al., 2015; Qin et al., 2021a; Bosman et al., 2022; Qin et al., 2024) with reionization likely driven by low-mass halos (e.g. Kuhlen & Faucher-Giguère, 2012; Qin et al., 2021b; Saxena et al., 2024; Mutch et al., 2024). However, the impact of Pop. III stars and mini-halos on the EoR is unclear. Pop. III stars are likely to be the dominant contribution to the total star formation rate density (SFRD) at and, if their IMF is more top-heavy than the present day one, Pop. III could significantly contribute to the heating and the ionization of the IGM which determines the evolution and shape of the 21cm signal (e.g. Qin et al., 2020; Gessey-Jones et al., 2022; Sartorio et al., 2023).
@@
The 21cm signal represents our most promising tool to put constraints on the thermal state of the IGM during the Cosmic Dawn and EoR. Even though no confirmed detection has been reported so far, the first upper limits on the 21cm power spectrum obtained with HERA phase I strongly disfavour cold reionization scenarios (HERA Collaboration et al. 2023). In the last few years, the impact of Pop. III stars on the 21cm signal has been studied using both analytical and semi-analytical models (e.g. Cohen et al., 2017; Chatterjee et al., 2020; Mebane et al., 2020). However these models either did not compute reionization (e.g. Magg et al. 2022; Hegde & Furlanetto 2023; Cruz et al. 2024a), focusing only on the absorption trough of the 21cm signal occurring at , or used a very simple analytical approach to compute reionization (e.g.Ventura et al. 2023). On the other hand, Cohen et al. (2017), Qin et al. (2021a) and Muñoz et al. (2022) used a simple analytical model for modeling Pop. III star formation but computed the reionization self-consistently.
@@
In this work, we overcome these challenges using a realistic Pop. III star formation and mini-halo model (Ventura et al. 2024) developed within the semi-analytical model meraxes designed to self-consistently couple galaxy formation and reionization. While in Ventura et al. (2024) we ran this model on a small (L = 10 h-1 cMpc) and high-resolution box, here we extend it to a significantly larger volume simulation (L = 210 h-1 cMpc) enabling the study of cosmic reionization. Since at such large volumes we cannot directly resolve mini-halos, we implemented scaling relations between the SFRD and the dark matter density field calibrated on the results from the small and high-resolution box discussed in Ventura et al. (2024). With this new model we are able to accurately follow the evolution of the radiative backgrounds relevant to the EoR and 21cm signal (X-rays, Lyman-, ionizing UV, Lyman-Werner) and to disentangle the contribution of Pop. III star formation to the 21cm global signal and power spectrum. Pop. III stars are expected to have a stronger impact at where they dominate star formation and ionization. Differently from previous works who explored the differences in the 21cm signal at Cosmic Dawn due to various Pop. III models, here we focus our attention on the residual signature of Pop. III on the 21cm power spectrum at where the sensitivity of SKA is expected to be significantly better and a detection is more plausible. To achieve this, it is crucial to model both Pop. III star formation and reionization in a self-consistent framework. This study allows us to assess under which conditions an early heating of the IGM from Pop. III stars leaves a detectable imprint on the 21cm power spectrum at .
@@
This paper is structured as follows. In Section 2 we give a brief overview of Pop. III star formation in meraxes. In Section 3 we present the scaling relation between the SFRD in mini-halos and the density field calibrated from the small and high-resolution box which is implemented in the large (210 h-1 cMpc)3 box. In Section 4 we discuss the impact of different Pop. III star formation models on the 21cm power spectrum and in Section 5 we make forecasts on the observability of these power spectra with SKA. Finally, we summarize our main results and conclusions in Section 6. Our simulations use the best-fit parameters from the Planck Collaboration et al. (2016): h = 0.6751, = 0.3121, = 0.0490, = 0.6879, = 0.8150, and ns = 0.9653.
@@ 2.1 Galaxy formation
In V24 the cooling prescriptions were updated to account for H2 cooling (the main cooling channel in mini-halos) and a more detailed metal enrichment model to keep track of the metallicity of each gas reservoir in a halo (which is crucial to distinguish between Pop. III and Pop. II star formation episodes). We also account for the effects of both baryon-dark matter streaming velocity and H2 photo-dissociation by the Lyman-Werner background which increases the minimum mass of a mini-halo capable of hosting stars (e.g. Schauer et al., 2021). Our model accounts for spatial variations only for the LW background, while for the relative velocity we assume a mean value throughout the entire box. For this work, we updated meraxes by adding the effect of H2 self-shielding which counteracts the H2 photo-dissociation, increasing the Pop. III SFRD by up to one order of magnitude at (see e.g. Feathers et al. 2024). We discuss the details of the implementation and impact of H2 self-shielding in meraxes in the Appendix A.
@@
The main free parameters that regulate the galaxy formation in meraxes are summarized in Table 1. The Pop. II related ones are taken from Balu et al. (2023a) where meraxes was run on a cosmological volume of L = 210 h-1 cMpc resolving all atomic cooling halos and calibrated in order to match the observed UV luminosity functions at (the agreement holds up to as shown in Qin et al. 2023) and the stellar mass functions at . Given the lack of observations of Pop. III stars, the Pop. III parameters are largely unconstrained. The fiducial values adopted in this work are taken from V24 and their values are suggested from hydrodynamical simulations (e.g. Chon et al., 2021).
@@ 2.2 Reionization and radiative backgrounds
where Tα is the colour temperature which we take equal to TK while xα and xc are the Lyman- and collisional coupling coefficients, respectively. These coefficients quantify the strength of the processes (resonant scattering of Lyman- photons, Wouthuysen 1952, and collisions with free electrons) that drive the spin temperature towards the kinetic temperature (when xα + xc 1, T TK otherwise TTγ). TK is sensitive to the adiabatic cooling and to all the processes able to heat up (or cool) the IGM with the most dominant coming from the X-ray emission. Hence in this work we will consider only the X-ray heating neglecting the other source of heating such as primordial magnetic fields (Minoda et al. 2019; Bera et al. 2020; Cruz et al. 2024b), Lyman- (Ciardi et al. 2010; Reis et al. 2021; Mittal & Kulkarni 2021), shocks (Furlanetto & Loeb 2004; Gnedin & Shaver 2004; Ma et al. 2021), cosmic rays (Bera et al. 2023), early accreting black holes (Mebane et al. 2020; Ventura et al. 2023) and decaying or annihilating dark matter (Liu & Slatyer 2018; Sun et al. 2023; Hou & Mack 2024; Facchinetti et al. 2024). Ultimately, the evolution of TS during the Cosmic Dawn and the EoR is mostly determined by the Lyman- (for the coupling between TS and TK) and X-ray flux.
@@
In this section, we present a novel approach that enables us to efficiently estimate the SFRD from mini-halos in a large box (that does not directly resolve these objects) using results from a small, high-resolution box. In Table 3 we summarize the key parameters for both the small (L10) and large (L210) box.
Label | Box side (cMpc) | Mass resolution (M⊙) | Pixel side |
L10 | 10 h-1 | 4.71 | 0.2 h-1 |
L210 | 210 h-1 | 3.16 | 0.2 h-1 |
@@ 3.1 Calibrating scaling relations from the L10 box
In Fig. 3 we show the redshift evolution of these parameters for 0.5 (black), 1.0 (grey), 1.5 (purple), 2.0 (red), 2.5 (green) and 3.0 (blue). exhibits an almost constant trend in redshift and a correlation with (higher results in higher ). This demonstrates that is mostly determined by the number of Pop. III star forming halos in a pixel, which is higher for more overdense regions. Since Pop. III star formation episodes in mini-halos are often the first episode of star formation experienced by a galaxy, it is not impacted by supernova feedback222Even though supernova feedback can be neglected, this is not true for the Lyman-Werner background that halts star formation in mini-halos. For this reason we add a further condition that if a pixel is irradiated by a LW flux J Jcrit above a critical threshold defined as M Mato (the atomic cooling threshold), (log10(SFR) = 0. Mato is the virial mass correspondent to a halo with a virial temperature T K. while Mcrit,MC is defined in Eq. 8. so the Pop. III SFR is almost constant at all redshift.
@@ 3.2 Applying scaling relations to the L210 box
Differently from what has been done by Hazlett et al. (2024) who calibrated a semi-analytical model to the Reinassance simulation in order to account for Pop. III star formation by adding the previous star formation history to each atomic cooling galaxy resolved in the simulation, the methodology described in this section allows us to estimate only the total SFRD occurring in mini-halos within a certain pixel of the simulation. In this work we focus on three models of Pop. III star formation by varying three main parameters: the star formation efficiency, the IMF and the specific Pop. III X-ray luminosity per unit star formation while all the other free parameters (e.g. the escape fraction) are fixed at the fiducial value (see Tables 1 and 2). We chose to focus only on these three parameters since these have the strongest impact on both the evolution of the Pop. III SFRD and the amount of UV and X-ray photons emitted. The values chosen for the specific Pop. III X-ray luminosity per unit star formation are similar to those found in Sartorio et al. (2023) for the different IMFs explored in their work. Hereafter, we analyze three different Pop. III models designed to have the minimum, intermediate and maximum impact from Pop. III star formation in mini-halos, each model is summarized in Table 4. The IMFs considered in this model are a Salpeter between 1 and 500 solar masses and a log-Normal IMF centered at 60 M⊙ (see V24 for more details). We highlight that in our extreme Pop. III model we enhance star formation efficiency, LX/SFR and the top-heaviness of the IMF at the same time. Each of these parameters has a different impact on the evolution of the 21cm signal (see Appendix B for a more detailed discussion of how each Pop. III parameter changes the evolution of the 21cm global signal and power spectrum).
@@
Extreme Pop. III | logE | 0.08 | |
High SFE | Salpeter | 0.08 | |
LogE | logE | 0.008 |
@@
We compute the sky-averaged 21cm global signal (see Fig. 10) without (black line) and with (grey, cyan and red line for weak, moderate and extreme Pop. III respectively) Pop. III star formation. As expected, introducing a Pop. III population with the same X-ray properties as Pop. II ones (i.e. weak Pop. III), simply shifts the absorption through to earlier epochs in virtue of the stronger coupling at higher- (see also Ventura et al. 2023; Hegde & Furlanetto 2023). However, if Pop. III stars have a stronger X-ray emission (i.e. moderate and extreme models) as suggested by Sartorio et al. (2023), the absorption signal is quickly suppressed turning into an emission signal as early as for the extreme Pop. III model and at 13 for the moderate Pop. III one. We note that a similar result has been found by a contemporaneous work by Gessey-Jones et al. (2025) who found an analogous variation in the timing () and depth ( mK) of the absorption trough when considering a stronger X-ray contribution from Pop. III stars (in their model the LX/SFR is self-consistently modeled from the IMF so that the difference between their Int-0 and Sal model is of 2 orders of magnitudes.)
@@
As shown in Fig. 11 earlier coupling and heating from Pop. III impacts the 21cm power spectrum both at the large and small scales. The models considered here produce 21cm signals that are different not only during the coupling and heating epoch (), but also at lower redshift when the reionization is in progress. The impact is stronger at smaller scales where models with a stronger heating exhibit a larger power spectrum at . This is of crucial importance as current and upcoming facilities are improving their observations at 10 (see more discussion in Section 5).
We compare our results with the ones in Muñoz et al. (2022) who studied how different Pop. III parameters impact the 21cm signal using 21cmFast. In the left panel of Fig. 11 the yellow and brown dashed lines are obtained by changing the specific Pop. III X-ray luminosity by a factor of 9. The global trends are similar with our models with stronger Pop. III X-ray emission (dashed yellow line and red/cyan lines) showing an earlier and weaker first peak compared to low X-ray models (dashed brown and grey lines). In their model, different Pop. III X-ray properties strongly impact the position and amplitude also of the second peak. In our models instead, the position of the second peak is only slightly anticipated in the strong X-ray models. This difference is likely due to the fact that our second peak occurs at much lower redshift () when most of the emission comes from Pop. II stars. Finally, differently from Muñoz et al. (2022) at simulations with a stronger Pop. III X-ray heating have a significantly larger power spectrum compared to weak X-ray models (a similar result has been found also by Gessey-Jones et al. 2025).
@@
This analysis shows that Pop. III star formation not only impacts the 21cm global signal during the cosmic dawn as previously assessed (e.g. Qin et al., 2020; Gessey-Jones et al., 2022; Ventura et al., 2023; Hegde & Furlanetto, 2023; Cruz et al., 2024a) but also that an early () heating of the IGM provided by this population leaves a strong signature on the power spectrum during the EoR (). The impact on the power spectrum is stronger for models that have a large Pop. III X-ray emissivity (i.e. moderate and extreme Pop. III models) while models with a milder X-ray emission (i.e. weak Pop. III) have a stronger effect on the global signal. Ultimately, this tells us that the power spectrum during the EoR can be used to disentangle different heating models and potentially constrain the properties of Pop. III stars. We highlight that here we considered only the X-ray emission from stellar remnants. While there might be other sources that significantly heat the IGM at (see discussion at end of Section 2.3), the other effects are likely to be either subdominant, still related to Pop. III stars (e.g. cosmic rays) or dominant only at the dark-ages (e.g. dark matter annihilation).
Finally, we note that in this work we did not include the effect of the X-ray feedback on Pop. III star formation. As noted by Ricotti (2016); Park et al. (2021), X-rays have both a positive and negative effect on Pop. III star formation as they heat the gas and increase the electron fraction. The heating makes gas accretion more difficult (hence delaying star formation) and free electrons promote the formation of H2 making the molecular cooling more efficient. In presence of a strong X-ray background this latter effect is dominant at (see e.g. Fig. 9 in Hegde & Furlanetto 2023). Hence, including the X-ray feedback would likely increase the Pop. III SFRD making the impact of Pop. III stars even stronger than what predicted here especially for the moderate and extreme models.